OOB connectivity remapping
Resolved Mar 24 at 10:30am GMT
We will be doing work on our out-of-band network to allow for better maintenance and to resolve some instability issues with our automatic services.
During this window some services ability to reboot and reinstall may be disrupted.
Upgrading uk-core switch
Resolved Mar 24 at 09:01am GMT
This maintenance has begun, we do not expect ongoing disruption of more than 5 minutes.
1 previous update
Resolved Mar 07 at 02:00pm GMT
During this timeframe the API might be rendered unresponsive for a brief period of time.
Emergency maintenance to update critical switching firmware
Resolved Mar 04 at 12:49am GMT
Service has been restored, we will provide affected customers with a more accurate rundown of the outage details as well as applicable compensation in the coming days.
4 previous updates
Metal deploy migration
Resolved Feb 21 at 01:30pm GMT
We are currently migrating some internal services to our new operations boxes in order to increase redundancy and reliability.
This migration was successfully completed at 13:53 GMT.
Emergency routing changes
Resolved Feb 17 at 03:30pm GMT
We will be performing a BGP session swap in London.
This should not be disruptive to clients.
There was a brief disruption to services in our London location at 23:20pm GMT this was due to a routing error with our upstream when we swapped our BGP session. This has been resolved and no further disruption is expected.